. A corporate note is bought on 5/3/2005 and matures to $4000 on 5/9/2005. On 

5/7/2005 this note is sold to a third party requiring 7%. Find the amount the 

third person paid for the corporate note. Find the actual interest the original 

holder lost on his investment.


1
Expert's answer
2021-04-29T18:10:39-0400

(a) 4280-4000= 280


(b) 280=4000*R/100*1/3

40R=840

R=21%

4000*21/100*1/2=420

420-280=140